10.6 HOW THE FEDERAL RESERVE CONTROLS THE INTEREST RATE LEARNING OBJECTIVE: Understand how the Fed can change the interest rate. 6.1 In the Republic of Doppelganger, the currency is the ditto. During 2015, the Treasury of Doppelganger sold bonds to finance the Doppelganger budget deficit. In all, the Treasury sold 80,000 ten-year bonds with a face value of 1,000 dittos each. The total deficit was 80 mil- lion dittos. The Doppelganger Central Bank reserve requirement was 16 percent and in the same year, the bank bought 10 million dittos' worth of outstanding bonds on the open market. All of the Doppelganger debt is held by either the private sector (the public) or the central bank. a. What is the combined effect of the Treasury sale and the central bank purchase on the total Doppelganger debt outstanding? On the debt held by the private sector b. What is the effect of the Treasury sale on the money supply in Doppelganger?
CHAP Assuming no leakage of reserves out of the bankin stem, what is the effect of the central bank purchase of sy bonds on the money supply?
Answer #1

SOLUTION:

1) Doppelganger outstanding debt will be a change in overall debt of 80,000,000 dittos (i.e. 80,000 shares @ $1,000) and, in specific the debt held by the private sector rises by 70,000,000 dittos because bank bought 10,000,000 rags worth

2) The Treasury sale will not impact the Doppelganger's money supply thus the result will be 0 dittos.

3) If we suppose that ther are no leakage of reserves out of the banking system, the effect of purchase of bonds on the money supply on of the Doppelganger Central Bank will be change of 62,500,000 dittos,

Multiplier = 1 /0.16 = 6.25

Thus result of 10 million shares bough back will be 6.25 * 10,000,000 = 62,500,000